suppose θ is polarising angle for a transparent medium and speed of light in that medium is v then according to brewester law

Grade:12th pass

1 Answers

Susmita
5 years ago
Light falls on a vaccum of refractive index 1 at an angle theta and it is refracted to a medium having refractive index n.
Refractive index n=c/v
where c velocity of light in vaccum and v velocity of light in medium.
According to brewsters law,
\theta = tan^{-1} (\frac{n}{1})=tan^{-1}(c/v)
Hope this helps.